### v1

# Implicit State Estimation via Video Replanning

Abstract
Video-based representations have gained prominence in planning and decision-making due to their ability to encode rich spatiotemporal dynamics and geometric relationships. These representations enable flexible and generalizable solutions for complex tasks such as object manipulation and navigation. However, existing video planning frameworks often struggle to adapt to failures at interaction time due to their inability to reason about uncertainties in partially observed environments. To overcome these limitations, we introduce a novel framework that integrates interaction-time data into the planning process. Our approach updates model parameters online and filters out previously failed plans during generation. This enables implicit state estimation, allowing the system to adapt dynamically without explicitly modeling unknown state variables. We evaluate our framework through extensive experiments on a new simulated manipulation benchmark, demonstrating its ability to improve replanning performance and advance the field of video-based decision-making.
1 Introduction
Learning from videos has gained significant traction in decision-making, as videos capture rich visual and dynamic information while aligning with how humans acquire knowledge. These properties make them a powerful medium for specifying tasks and learning diverse skills across contexts. Recent work has shown the effectiveness of video-based frameworks in enabling robots to learn behaviors such as object manipulation Li et al. (2024) and navigation Zhang et al. (2024), highlighting the value of video as a flexible and expressive representation.
This paper focuses on video as a planning representation. Given a goal and current observation, video planning systems generate imagined task executions and convert them into robot actions. Unlike symbolic or latent representations, videos naturally encode both perceptual and action information and generalize across tasks and environments. Prior works Chang et al. (2020); Du et al. (2024a; b) leverage these properties to train universal agents using video-based predictions.
Despite promising results, existing video planning frameworks suffer from a crucial limitation: they lack mechanisms to integrate past interactions with the environment and cannot effectively reason about uncertainty due to partial observability. Consider the task of opening a door without knowing whether it should be pushed or pulled. If pushing fails, a human will naturally infer that pulling is the correct action and adjust accordingly. In contrast, current video planning frameworks simply generate a new plan, disregarding the knowledge gained from prior attempts. This inability to incorporate feedback from interactions significantly limits their effectiveness, particularly in unstructured environments where uncertainty is inherent.
A common solution introduces explicit belief models trained in simulation to infer hidden parameters Memmel et al. (2024); Qi et al. (2023). Yet these methods require prior knowledge of relevant parameters and may struggle when interaction data fails to disambiguate uncertainties. In contrast, humans can effectively resolve these challenges by drawing on past experiences and employing adaptive trial-and-error strategies when prior knowledge is lacking. In the context of video planning, failed interactions are not merely setbacks — they offer critical information about the environment and system parameters.
Figure 1: Implicit state estimation via video replanning. Previous video planning methods rely on the first frame (in red). We propose to incorporate past failures (in blue) to improve future plans.
Our goal is to develop a video planning framework that effectively encodes and retrieves interaction data, adapting plans dynamically based on past experiences—without requiring additional simulation data or prior knowledge of the parameters. We propose a novel problem formulation that formalizes the challenge of generating video plans while implicitly incorporating past interactions, as illustrated in Figure 1. To address this, we introduce implicit state estimation (ISE), a framework that integrates interaction-time data into the planning process without explicitly defined system parameters. Instead of relying on hand-specified beliefs, our approach enables rapid adaptation by optimizing a subset of model parameters that implicitly encode unobservable state features. ISE also incorporates a plan-rejection mechanism that filters out previously failed plans, preventing over-reliance on incorrect inferences and encouraging exploration. By adopting a video planning framework, our approach leverages the latest advancements in generative models, paving the way for future applications such as out-of-distribution detection.
To validate the effectiveness of our framework, we introduce the Meta-World System Identification Benchmark, a new dataset designed to evaluate online adaptation in decision-making across diverse robotic tasks with unknown system parameters. Experiments show that our method allows for rapid adaptation without explicit system identification. Compared to baselines, including RL methods and other methods for video planning, our approach significantly reduces replanning failures and generates more accurate video plans in dynamic, uncertain environments.
2 Related Works
Video planning.
Videos have emerged as a popular medium for planning in the physical world. Yang et al. (2024b). For planning actions, people have explored using video to train dynamics model Yang et al. (2024a); Ajay et al. (2023); Zhou et al. (2024); Qin et al. (2024) or reward model Ma et al. (2023); Chen et al. (2021a); Escontrela et al. (2023). Another popular approach involves directly predicting video sequence Du et al. (2024a); Ko et al. (2024); Black et al. (2024) or derived representations such as 3D-correspondence or point tracks to guide downstream policy Yuan et al. (2024); Wen et al. (2024); Bharadhwaj et al. (2025). However, most of these frameworks do not consider environmental uncertainty, which limits their ability to efficiently adapt and replan, often resulting in suboptimal performance in such environments.
System identification.
System identification focuses on learning system dynamics and estimating parameters using experimental data, with early work on input selection and the Fisher Information Matrix Schön et al. (2011); Ljung (1998); Menda et al. (2020). This research laid the groundwork for active identification strategies, which aim to maximize the amount of information gained from system inputs, a key element in classical system identification Hjalmarsson et al. (1996); Lindqvist & Hjalmarsson (2001); Gerencser & Hjalmarsson (2005). Recent advances in the field have applied these methods to real-world systems, such as identifying physical parameters Xu et al. (2019); Kumar et al. (2019); Mavrakis et al. (2020); Gao et al. (2020; 2022); Memmel et al. (2024). Concurrently, Li et al. (2025) has explored verifier-based frameworks that reject infeasible action proposals based on historical input, enabling online adaptation. Our work can be seen as bridging classical system identification and modern video planning method by injecting the ability to implicitly infer physical parameters from past video data to the model.
Imitation learning.
Imitation learning Zare et al. (2024) is a research field of robot learning, aiming to train an agent to behave like an expert, given the dataset of the interaction of the expert and the environment. Imitation learning has been proven to be an effective and data-efficient approach for robot learning Ho & Ermon (2016); Wang et al. (2017); Lee et al. (2021); Lai et al. (2024); Huang et al. (2024). Behavioral cloning (BC; Torabi et al., 2018) formulates imitating an expert as a supervised learning problem and achieves success on various tasks Florence et al. (2021); Chen et al. (2024); Shridhar et al. (2022); Chi et al. (2024); Brohan et al. (2023); Yeh et al. (2025). Imitation learning uses successful expert demonstration, while we aim to leverage failed interactions with environments.
Offline reinforcement learning.
Offline RL assumes datasets composed of state-action-reward sequences Levine et al. (2020); Figueiredo Prudencio et al. (2024). In contrast to online RL, during training, the agent tries to learn to perform the task by leveraging the dataset, without any interaction to the environment. The paradigm of offline RL has existed for dacades. However, with the rise of deep learning, most modern offline RL algorithms utilize deep neural networks to express the value functions and policies Kidambi et al. (2020); Yu et al. (2020); Kumar et al. (2020b); Yu et al. (2021); Yang et al. (2022). In addition, with the development of large language model and transformer, there are also some research about using transformer for reinforcement learning Chen et al. (2021b); Furuta et al. (2022). Unlike offline RL, our framework do not assume dense reward signals.
3 Problem Formulation
We consider a setting where certain system parameters $\theta$, such as mass, friction, or utility — essential for solving the task — are unknown. Unlike existing work for explicit state estimation, we do not assume access to the parameterization of $\theta$. We only assume that the underlying $\theta$ can be inferred from interaction data. For example, in a bar-pushing task, the center of mass can be implicitly inferred from an interaction video, regardless of whether the push is successful.
We aim to design a system capable of rapidly solving tasks by planning efficiently and leveraging interactions in the environment. For each environment, we assume access to an experience dataset $\mathcal{D}$, consisting of data tuples $d_{i}=(v_{i},o_{i},s_{i})\in\mathcal{D}$, where $v_{i}$ represents an interaction video, $o_{i}$ denotes the object ID, and $s_{i}$ is a binary success indicator. The dataset contains both successful interactions where $s_{i}=1$ and failed interactions where $s_{i}=0$. Our work focuses on evaluating whether the selected plans can guide successful action sequences. We complement our focus on plan evaluation by using a fixed action prediction module to convert videos into actions for execution.
4 Approach
Our world is full of uncertainty. For example, when trying to move a heavy piece of furniture, you often don’t know its exact weight, how slippery the floor might be, or whether its legs will flex under pressure. As humans, we actively make new plans and use interaction trials to adapt: infer hidden parameters and narrow the space of possible dynamics.
Our framework follows this intuition. In the video planning framework, we propose to leverage prior interaction videos to refine the agent’s belief about the environment through trial-and-error and video-based planning. Shown in Figure 2, given the current scene and past interactions, it retrieves and updates a latent state embedding that encodes a hypothesis of the hidden parameters, then generates a video plan conditioned on this embedding.
Concretely, our video replanning framework maintains two test-time buffers: one for failed interactions and another for failed plans. At each round, it estimates hidden parameters from failed interactions, then generates new candidate plans guided by this estimation (Section 4.1). A rejection strategy (Section 4.2) selects the plan most different from past failures. The chosen plan is then converted into executable actions via an Action Module (Section 4.3), which can be instantiated with inverse dynamics models, goal-conditioned policies, or trajectory-tracking controllers. In our implementation, we adopt the training-free point-tracking method from AVDC Ko et al. (2024). For extensive details of each component, please refer to Appendix C.
Figure 2: Framework overview. (a) Generate and select plans The Video Plan Generator first generates a set of candidate video plans conditioned on first-frame $f_{0}$ and state embedding $e$. The Rejection Module then selects the plan that is least similar to the plans in the Failed Plan Buffer $B_{P}$. (b) Execute plan and update buffers. The Action Module interacts with the environment by following the selected plan $v^{p*}$. If the interaction trial is not successful, we update the Failed Plan Buffer $B_{P}$ and the Failed Interaction Buffer $B_{I}$. (c) Update belief. The state embedding $e$ is updated using the Retrieval Module based on the Failed Interaction Buffer $B_{I}$. Then, we generate the next batch of video plans according to the updated belief.
4.1 Video-based Planning with State Embeddings
To plan effectively in environments with unknown dynamics, our system constructs a latent representation of hidden parameters, such as mass or friction, through state embeddings. These embeddings act as hypotheses about the environment configuration and serve as conditioning signals for generating scenario-specific rollouts. By combining prior retrieval with test-time refinement, our framework maintains state embeddings as a flexible and adaptive representation of the environment.
Representing state embeddings.
Interaction videos from prior executions with the same object are encoded using a context encoder $E$, producing video embeddings $e^{v}_{j}=E(v_{j})$. To ensure consistency across multiple trials of the same object, we preprocess the dataset $\mathcal{D}$ by grouping entries by object ID and selecting a successful trial as the canonical embedding $e^{o}_{j}$. The choice of encoder $E$ defines the latent feature space; we experiment with pretrained vision encoders such as CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2024), which yield semantically structured spaces where different environment modes can be distinguished. Please refer to Section D.1 for visualization of such latent space and Section D.2 for details of dataset $\mathcal{D}$ used in our experiments.
Figure 3: Video Plan Generator and Identification Module. We implement the Video Plan Generator and the Identification Module as two video generative models with shared parameters, switched with a binary trigger is_planning. The Video Plan Generator only learns from successful videos in the experience dataset $D$, while the Identification Module is trained with all videos in $D$, including the failures, enabling the state-embedding refining process.
Retrieval and refinement.
At inference time, the planner requires a state embedding aligned with the current interaction. To obtain such an embedding, we retrieve candidates from prior interactions by comparing the features of the current video against stored embeddings. Instead of deterministically selecting the closest match, we treat the distances as similarity scores, normalize them with a softmax function, and sample a state embedding according to the resulting probability distribution. This stochastic retrieval helps the planner remain robust when multiple past interactions are plausible. However, retrieval alone can still misalign with the ongoing interaction. To address this, we introduce an optimization-based refinement step illustrated in Figure 3: a generative identification module, which shares parameters with the Video Plan Generator, is trained to produce both successful and unsuccessful outcomes. At inference time, we freeze the module’s weights and optimize the embedding to minimize the discrepancy between its generated rollouts and the observed interaction using a standard denoising diffusion loss. This refinement progressively adjusts the retrieved embedding toward the true environment dynamics, providing a more faithful basis for planning.
Video plan generation.
Once a state embedding is obtained, the Video Plan Generator uses it to predict a sequence of plausible futures conditioned on the current observation. We adopt a diffusion-based planner following AVDC Ko et al. (2024), but replace language-based conditioning with state embeddings. Specifically, the embedding is projected into the token space of CLIP-Text and appended as an additional conditioning token to the 3D U-Net backbone. The generator is trained end-to-end with a denoising diffusion objective, which encourages realistic rollouts under the hypothesized environment. In practice, we generate M=7M{=}7 future frames at $128{\times}128$ resolution and concatenate them with the initial observation to form an 8-frame video plan. These plans capture not only the high-level task intent but also environment-specific details encoded in the embedding, making them a suitable substrate for downstream rejection and action execution.
4.2 Rejection-based Replanning
While the retrieval and refinement procedure introduced in Section 4.1 provides a plausible hypothesis of the environment, it is rarely perfect, especially with limited interaction data or in previously unseen situations. To improve robustness, we introduce a rejection module that actively explores alternatives, encouraging the planner to avoid repeating failed strategies.
Candidate plan sampling.
At each planning round, instead of generating a single rollout, the Video Plan Generator produces $N$ candidate plans $\{v^{p}_{1},v^{p}_{2},\dots,v^{p}_{N}\}$ by conditioning on $N$ independently sampled state embeddings from the Retrieval Module. Because each sampled embedding represents a different hypothesis of the environment parameters, the resulting video plans can differ substantially, covering a range of possible outcomes.
Rejection strategy.
To ensure the planner does not simply repeat failed behavior, we maintain a buffer $\mathcal{F}$ of previously failed plans. For each candidate plan $v^{p}_{i}$, we compute its distance to the closest failed plan in $\mathcal{F}$:
$d(v^{p}_{i},\mathcal{F})=\min_{F\in\mathcal{F}}K(v^{p}_{i},F),$
where $K$ denotes the distance function. Empirically, we found that using an $L2$ distance directly on the pixel space of video frames works well. The rejection module then selects the plan with the largest distance to past failures,
$v^{p*}=\arg\max_{v^{p}_{i}}d(v^{p}_{i},\mathcal{F}),$
which is subsequently passed to the Action Module for execution. This procedure encourages the planner to explore novel strategies rather than repeating unsuccessful ones, allowing the system to refine its behavior through active interaction.
4.3 Action Module
The final step is to convert a selected video plan into executable low-level actions for the robot. We follow the idea of dense object- or wrist-tracking from AVDC Ko et al. (2024): the predicted video frames are used to track either the object of interest or, in tasks such as bar picking, the robot’s wrist when object tracking is insufficient. The resulting trajectory is then translated into control commands through simple heuristic policies such as grasp strategies or trajectory-following controllers. This training-free approach enables direct execution of video plans without requiring an additional learned policy.
5 Experiment
5.1 Meta-World system identification benchmark
Existing manipulation benchmarks primarily emphasize single-episode task completion while assuming full observability, overlooking the rich information available through interaction. Consequently, they fail to capture challenges posed by environmental variation, such as determining whether a door should be pushed or pulled, that humans frequently encounter in real world and can solve with ease. Therefore, in this work, we propose the Meta-World System Identification Benchmark to evaluate online adaptation under unknown system parameters $\theta$. It combines an offline experience dataset with environments exhibiting hidden parameters that agents must infer through interaction. Table 1 shows the set of 5 tasks, including object manipulation (e.g., push bar, pick bar) and puzzle solving (e.g., turn faucet), test adaptability to randomized centers of mass, uncertain interaction modes, and external forces.
Push bar. The objective of this task is to push a bar toward a designated target with the center of mass of the bar randomized.
Table 1: Meta-World System Identification Benchmark. Each task involves interaction-based inference of system parameters $\theta$, which are either continuous or discrete with varying mode counts.
Task Identification Type
Push Bar Continuous
Pick Bar Continuous
Slide Brick Continuous
Open Box Discrete, 2 modes
Turn Faucet Discrete, 2 modes
Pick bar. This task requires the robot to grasp a bar and transport it to a specified target. Similarly, the center of mass of the bar is randomly assigned.
Slide brick. This task consists of two distinct stages. In the first stage, the robot must push a brick up an inclined surface. In the second stage, the brick slides freely down the slope. The objective is to control the initial push such that the brick comes to rest within a predefined target region on the slope.
Open box. The objective of this task is to open a box. The box cover can be manipulated in one of two modes: it can either be lifted or slid open.
Turn faucet. In this task, the robot must rotate a faucet. The faucet’s direction of rotation is either clockwise or counterclockwise.
In the above five tasks, the system parameters cannot be inferred without physical interaction. The details of the tasks are provided in Appendix B.
5.2 Evaluation metrics
Our study focuses on two key aspects. First, we hypothesize that our replanning mechanism can reduce the number of replans required for successful execution. To quantitatively measure the rate of adaptation, we report the average number of replans needed per successful attempt in our main experiment. Specifically, the agent must repeatedly replan until it either succeeds or reaches a predefined maximum number of replan trials, $M$. For all tasks, we set $M=14$.
Second, instead of only evaluating task success, we also assess the effectiveness of the replanning mechanism by comparing the replanned videos to groundtruth interaction videos. To quantify this similarity, we use standard video comparison metrics: PSNR, SSIM, and LPIPS Zhang et al. (2018).
Figure 4: Qualitative results of adapted video plans. We show qualitative examples of adaptation trials for each task in the Meta-World Identification Benchmark to illustrate the effectiveness of the proposed method. The framework generates new plans based on previously failed interaction and video plans.
5.3 Implementation and baselines
We evaluate our framework on the 5 tasks in the Meta-World System Identification Benchmark. For our model, the Video Plan Generator is trained on the videos in the offline experience dataset as described in 4.1. Each task is evaluated over 400 trials. We compare our approach with the original AVDC implementation for video planning which does not online adapt, variants of our method that use different context encoders $E$, and other reinforcement learning baselines.
  • AVDC (Ko et al., 2024). The baseline with only the Action Module, i.e., without the Retrieval Module and Rejection Module introduced in this paper.
  • Ours. Our proposed method with $E$ based on DINOv2 feature as described in 4.1.
  • Ours (Refine). A variant of our method with refining state embeddings from scratch as mentioned in 4.1.
Our problem formulation for state estimation can be viewed as a single-step multi-task reinforcement learning (RL) problem, where the success signal $s_{i}$ from the dataset serves as the reward, by additionally assuming privileged access to action labels (i.e., ground-truth system parameters). We also implement such explicit state estimation baselines with popular RL algorithms to calibrate the difficulty of the tasks.
  • BC uses a state embedding-conditioned behavior cloning framework where the Policy takes in state and state embedding directly. The state embedding is obtained with $E$ based on DINOv2.
  • CQL (Kumar et al., 2020a) is a state embedding-conditioned actor-critic framework. The Q-function takes in state, action, and state embedding, while the Policy takes in state and state embedding, which is obtained with $E$ based on DINOv2.
Table 2: Number of replans until success. Video planning methods, AVDC and ours, outperform baselines in imitation learning (BC) and offline RL (CQL) on most tasks. Our method and its variant consistently outperform the video planning baseline. The error terms show the standard error of the mean value. The reported error terms represent the standard error of the mean across 400 trials. For the CQL baseline, performance is averaged over three independently initialized and trained policies for more robust evaluation.
Method Push Bar Pick Bar Slide Brick Open Box Turn Faucet All (Normalized)
[…]
5.4 Results
The replanning efficiency of all methods is presented in Table 2. Our method and its variants consistently outperform baselines in video planning (AVDC), imitation learning (BC), and offline RL (CQL). AVDC performs better than BC and CQL in terms of trial efficiency, demonstrating the benefits of video-based planning. BC and CQL struggle significantly on these tasks and often fail to complete them, frequently repeating similar predictions—suggesting overfitting or limited flexibility. Overall, our approach achieves the strongest performance, showing its broad applicability across both discrete and continuous parameter settings.
Table 3: Evaluation on replanned videos. We quantitatively assess the accuracy of the generated video by measuring the similarity between the framework’s output plan and the ground-truth plan produced by the GT scripted policy.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
AVDC 19.426 0.700 0.271
Ours (R3M) 19.632 0.704 0.261
Ours (CLIP) 19.584 0.703 0.262
Ours (DINOv2) 19.817 0.708 0.255
Our variants. Among the variations of our method, using the Rejection Module (Ours) alone to generate state embeddings achieves the best overall performance. In contrast, the refinement process introduced in Section 4.1 alone, Ours (Refine), yields similar or slightly worse performance but does not assume data availability at test time, i.e., do not require access to the dataset used to train the diffusion model. Additionally, its formulation as a learned model offers the potential for improved generalization. Finally, we show qualitative results for adaptations in Figure 4.
Video plan evaluation. In Table 3, we report the evaluations of video plans generated by the baseline (AVDC) and our methods with various context encoders $E$, including R3M Nair et al. (2022), CLIP Radford et al. (2021), and DINOv2 Oquab et al. (2024). Our method produces significantly higher-quality videos compared to AVDC. Our method with DINOv2 outperforms its counterpart using R3M or CLIP, validating our design choices.
Figure 5: Rejection and retrieval modules. Both the Rejection and Retrieval Modules improve the performance from AVDC (Aciton Module only) and combining them (ours) yields the best performance. In this figure, the numbers of replans are normalized aggregated among the five tasks.
Figure 6: Rejection module distance. We compare using raw-pixel distance and DINOv2 embedding distance to reject and the former yields better performance. The numbers of replans are normalized and aggregated among the five tasks.
Figure 7: Number of generated video plans. We experiment with generating different numbers of video plans $n\in[1,5]$ and set $n=2$. In this figure, the numbers of replans are normalized and aggregated among the five tasks.
5.5 Ablation studies
Rejection and retrieval modules.
To investigate the effectiveness of the Rejection and Retrieval Modules, we compare AVDC (Action Module only), AVDC+Rejection, AVDC+Retrieval, and our method (with both Rejection and Retrieval). Notably, the AVDC+Rejection baseline is conceptually similar to a recent work Li et al. (2025), which also uses a generator to sample candidate plans and a separate module to select the most viable one given previous interaction attempts. The result in Figure 7 shows the normalized numbers of replans across all the tasks and individual task performance is shown in Appendix D. Both AVDC+Rejection and AVDC+Retrieval outperform AVDC, justifying the effectiveness of the Rejection and Retrieval modules. Our full method (with both the Rejection and Retrieval modules) achieves the best performance, validating the combinations.
Rejection module distance metric.
We compare pixel-wise distance and DINOv2-based similarity for selecting video plan candidates, as described in Section 4.2. Figure 7 shows that pixel-wise distance performs better, likely because DINOv2 struggles to distinguish these plans due to the nature of its pretraining data.
Number of generated video plans.
To study how the number of video plans generated to be rejected affects the performance, we vary the number from 1 to 5 and show the result in Figure 7. The result aggregates the performance across all the tasks. From $n=1$ (AVDC+Retrieval) to $n=2$ (ours), around $10\%$ of replans are saved. However, from $n=2$ to $n=5$, the numbers of replans until success are near the same. An increase in video plan candidates generated by the video generator results in an increase in the computational cost. Therefore, since the performances of $n\in\{2,3,4,5\}$ are comparable, we set $n=2$.
Rejection module distance metric.
To study the impact of the distance function used in the Rejection Module, we evaluated several alternatives. The results show that raw-pixel distance consistently achieves superior or comparable performance relative to the DINOv2 baseline. Based on this finding, we adopt raw-pixel distance as the default choice of $K$ for the Rejection Module.
Table 4: Rejection module distance metric. We compare the performance of using different distance functions for the Rejection Module.
Metric Push Bar Pick Bar Slide Brick Open Box Turn Faucet All (Normalized)
[…]
5.6 Evaluation on real world data
To demonstrate the method’s ability to plan efficiently, we evaluated the video-planning framework on a real-world door-opening task. The door can be either pushed or pulled open, which can’t be distinguished from the appearance and must be revealed through interaction. The setup is shown in Figure 8.
Figure 8: Real-world setup. A WidowX arm attempts to open a door that can either be pulled or pushed open.
We collected 20 real-world videos (10 successes, 10 failures) to train the video model. We address the challenge of training on such a small dataset by adopting a joint training with simulation data: 40% of samples came from real-world data and 60% from the Meta-World System Identification Benchmark. To reduce overfitting to the first frames, which often leads to monotonic plans, we injected Gaussian noise into the first-frames during inference. Furthermore, we set $n=4$ to promote diverse plan generation.
For simplicity, we evaluate the single-episode replanning success rate, conditioned on an initial failure—different from the main evaluation metric. A perfect random policy, which assumes knowledge of the correct modes (e.g., push or pull) but does not use past interaction, would succeed with 50% probability. Our method achieved 15/20 successes, compared to 8/20 for a naive video planner without using past information, demonstrating the effectiveness of our pipeline.
Figure 9: Real-world qualitative results. Given the first frame (a), the robot first attempts to push the door (b). Then, it gets stuck and fails to open the door (b). Finally, it generates a plan that pushes the door instead (c) by leveraging information from previous failure.
Figure 9 shows the qualitative rollout for our real-world experiment. We followed most of the hyperparameters as used in the Meta-World System Identification Benchmark. Additionally, we changed $n=4$ and injected Gaussian noise of standard deviation 0.35 into first-frames with pixel values normalized to $[0,1]$ to mitigate the first-frame bias issue.
6 Conclusion
We presented a video planning framework that adapts to interaction-time uncertainties by implicitly encoding past experiences into the planning process. Our method filters failed plans, updates models online, and achieves robust adaptation without explicit system identification. Experiments on the Meta-World System Identification Benchmark confirm improved robustness over video-based and RL baselines, highlighting the potential of video as a foundation for adaptive decision-making in uncertain environments.
Acknolwemgent
This work was supported by the Industrial Technology Research Institute, Taiwan, and the National Science and Technology Council, Taiwan (114-2628-E-002 -021-). Shao-Hua Sun was supported by the Yushan Fellow Program by the Ministry of Education, Taiwan.
Reproducibility Statement
We are committed to reproducibility and will release the full codebase for reproducing the experiments introduced in this work. Our experiments were conducted on standard GPUs (V100/3090) with moderate training times, and we include full details in the appendix to facilitate independent verification.
Appendix
Appendix A Supplementary materials
A.1 Website and Code
We present extensive qualitative results and release our code in our website.
Appendix B Benchmark Details of Simulator Tasks
Push Bar
Pick Bar
Slide Brick
Open Box
Turn Faucet
Figure 10: Settings of five task environments.
Figure 11: Push Bar. The above video shows a failed trial, because the deflection of the bar is too large so that the bar cannot reach the target.
Figure 12: Pick Bar. The above video shows a failed trial, because the bar fell down.
Figure 13: Slide Brick. The above video shows a failed trial, because the brick didn’t slide down to the gray band.
Figure 14: Open Box. The above video shows a failed trial, becuase the robot try to grasp the handle up. However, the right way to open the box is to push the handle horizontally.
Figure 15: Turn Faucet. The above video shows a failed trial, because the robot try to turn the faucet clockwise. However, the right way is to turn the faucet counterclockwise.
Figure 16: Visualization of simulation tasks.
The environments of the five tasks are shown in Figure 10.
B.1 Push Bar
The objective of this task is to push a bar towards the target. Physically, the bar is combined with a mass point and a mass-free bar. Therefore, if the robot does not push the bar exact at the center of mass, torque is imposed on the bar, resulting in the deflection of the bar.
The bar has a length of 0.4 meters. In the pretraining stage the video generation model, the distances between the center of mass and geometric center of the bar are shown in Table 5.
Table 5: Center of mass positions. The distances between the center of mass and the geometric center of the bar in the pretraining stage of the video generation model. The sign represents the direction.
-0.18 -0.165 -0.15 -0.135 -0.12
-0.105 -0.09 -0.075 -0.06 -0.05
-0.03 -0.015 0.0 0.015 0.03
0.045 0.06 0.075 0.09 0.105
0.12 0.135 0.15 0.18
The visualization of this task is shown in Figure 16.
B.2 Pick Bar
The objective of this task is to grasp a bar towards the target. Physically, like push-bar mentioned above, the bar is combined with a mass point and a mass-free bar. Therefore, if the robot does not grasp the bar exact at the center of the mass, torque is imposed on the bar, resulting in the tilt of the bar.
The properties of the bar in this task are the same as the bar in push-bar. And the distances between the center of mass and geometric center of the bar in the pretraining stage are also the same as above.
The visualization of this task is shown in Figure 16.
B.3 Slide Brick
There are two stages in this task. In the first stage, the robot pushes a brick up an inclined surface. In the second stage, the brick slides freely down the slope. The objective is to push the brick at a proper height so that the brick can slide down and stop at the gray band.
Note that the wooden slope has a relatively low friction coefficient and that the red slope has a relatively high friction coefficient. Therefore, the brick can slide down and accelerate in the wooden slope and decelerate in the red slope. The friction coefficient of the wooden slope is fixed and is set to 0, while the friction coefficient of the red slope is various, and those in the pretraining stage are shown in Table 6.
The visualization of this task is shown in Figure 16.
Table 6: Friction coefficients. The friction coefficients of the red slope in the pretraining stage of the video generation model.
0.24 0.25 0.26 0.27 0.28 0.3 0.32
0.34 0.35 0.36 0.38 0.39 0.4
B.4 Open Box
The objective of this task is to open a box whose lid operates in one of two possible ways: it either needs to be lifted or slid open. The specific mode is chosen at random and cannot be identified visually, so the robot must physically interact with the box and adjust its strategy based on the observed response.
The visualization of the task is shown in Figure 16.
B.5 Turn Faucet
The objective of this task is to rotate a faucet. The direction of rotation—clockwise or counterclockwise—is randomly assigned and cannot be determined in advance through observation. The robot must interact with the faucet to identify the correct direction and execute the appropriate motion.
The visualization of the task is shown in Figure 16.
Appendix C Method Details
C.1 Video plan generator and retrieval module
To generate meaningful plans in environments with unknown dynamics, our system first hypothesizes the hidden parameters of the environment via the retrieval and refinement of a latent state embedding, which serves as a proxy for the current environment configuration. We first discuss our Video Plan Generator and Retrieval Module, which generate candidate plans based on observed images and state embeddings. The Video Plan Generator predicts future frames conditioned on the current observation and, when available, the state embedding. For new objects without prior interactions, the generator operates solely on the observed image by conditioning on a learned “null” state embedding.
The Retrieval Module extracts state embeddings from prior interactions by encoding past videos and selecting a representative embedding for each object ID. At inference time, it encodes the current interaction video, computes distances to past embeddings from dataset $\mathcal{D}$, and uses a softmax function to sample the most relevant state embedding. In addition, we use an additional identification model to continuously refine the selected state embedding to best fit current interaction videos.
C.1.1 State embedding
To represent the hidden system parameters, we encode interaction videos from previous action executions with the same object using a context encoder $E$. The video embedding for each video $v_{j}$ is obtained as $e^{v}_{j}=E(v_{j})$. In the dataset $\mathcal{D}$, multiple data tuples may share the same object ID $o_{j}$. Naively encoding each video independently would result in multiple, different embeddings for the same object. To address this, we preprocess the dataset by grouping all entries with the same object ID and selecting the encoding from one of the successful executions (where $s_{j}=1$) to represent the object, resulting in a state embedding $e^{o}_{j}$.
C.1.2 Video plan generator
The Video Plan Generator is used to generate a video containing $M$ future frames conditioned on the current observation and a state embedding. We follow the implementation of AVDC’s video planner, which is a first-frame conditioned video diffusion model utilizing a 3D U-Net architecture. The original implementation conditions on task features were obtained from encoding natural language task descriptions with text encoders such as CLIP-Text, rather than state embeddings. To adapt the method to our needs, we project the state embedding into the hidden dimension of the language features and treat the projected embedding as an additional language token. The parameters of the projection layer are trained end-to-end along with the video planner, which is optimized using a standard denoising diffusion loss. For all experiments in this paper, we choose the resolution of generated frames to be 128x128 and $M=7$. Each time we invoke the Video Plan Generator, $M=7$ future frames are generated, we then concatenate them with the first frame, resulting in a $M+1=8$ frames video plan.
Table 7: Hyperparameters. Comparison of configuration parameters for the Meta-World benchmark.
   num_parameters    125M
   diffusion_resolution    (32, 32)
   target_resolution    (128, 128)
   base_channels    128
   num_res_block    2
   attention_resolutions    (2, 4, 8)
   channel_mult    (1, 2, 3, 4)
   batch_size    128
   training_timesteps    12k
   denoising_timestep    100
   sampling_timestep    25
Our video diffusion model adopts a U-Net architecture following (Ko et al., 2024). The original implementation conditions on the first frame and the task name tokens, where the task tokens are encoded using CLIP-Text, to predict future frames. Building on this framework, we additionally condition the video generation process on past interaction videos. Specifically, we encode the past interaction video as described in Section C.1.5, and concatenate the resulting embedding with the encoded task tokens after projecting it to match the CLIP-Text output dimensionality (512 channels) using a trainable linear layer.
To improve computational efficiency, the video diffusion model generates a low-resolution video plan through the denoising process. This low-resolution plan is then upsampled to the target resolution using a separate super-resolution model that predicts residuals over the upsampled frames. The diffusion model is trained using the v-parameterization approach with a learning rate of 1e-4. Key hyperparameters for the video diffusion model are summarized in Table 7.
C.1.3 Retrieval module
During inference, the Video Plan Generator needs to take in a state embedding extracted from prior interactions. To ensure that the method is general across different encoders $E$, we first normalize and align the dimensionality of the feature space constructed by video features $e^{v}_{j}$ by applying PCA to all extracted video features $e^{v}_{j}$, yielding a PCA projection $P$ and PCA-transformed features $e^{vp}_{j}=P(e^{v}_{j})$ and $e^{op}_{j}=P(e^{o}_{j})$.
Given the interaction video from a prior execution, $v^{e}$, we first encode it using the encoder $E$ and apply the PCA transformation, resulting in $e^{ep}=P(E(v^{e}))$. We then compute the distance between $e^{ep}$ and the video encodings from $\mathcal{D}$ as $logits_{j}=-dist_{j}=-K(e^{ep},e^{vp}_{j})$, where $K$ represents the distance function. For the choice of $K$, we used Cosine Distance for CLIP and DINO-based methods and L2 Distance for other methods. After calculating the logits, we compute the sampling probability of each output embedding by passing the logits through a softmax function: $p_{j}=\text{softmax}(logits_{j}/\tau)$, where $\tau$ is the temperature parameter. Finally, we sample the output state embedding $e=e^{op}_{k}$ according to the resulting probability distribution $k\sim\text{Categorical}(p_{1},p_{2},\dots,p_{n})$. While such retrieval-based state estimation formulation necessitates the data being available during test time, it enables training-free state estimation and guarantees that the Video Plan Generator never accepts out-of-distribution state embeddings during inference time.
C.1.4 Refining state embedding
In addition to using retrieval to find a relevant state embedding for a video, we also use an optimization-based approach to find a refined state embedding corresponding to a specific interaction video. Specifically, we introduce another generative model, referred to as the Identification model or the ID module, trained with the parameters shared with the Video Plan Generator, as shown in Figure 3. The ID module is trained to generate all possible interaction outcomes, including both successful and unsuccessful ones.
During replanning, we freeze its parameter and only optimize the state embedding to maximize the probability of generating the given interaction video.
$e=\arg\min_{e}MSE(v^{e},v)$
Here, $e$ is the optimized state embedding, $v^{e}$ is the interaction video obtained at test-time, and $v$ is the videos sampled from the ID module conditioned on $e$.
Since the ID module and the Video Plan Generator share the same embedding space for state embeddings, the optimized state embedding can be used directly to generate video plans. In the experimental section of the paper, we analyze the effect of this refinement procedure for state estimation when 1) this approach optimizes state embedding from randomly initialized vectors drawn from unit variance Gaussian distribution, 2) the state embedding is set to the output of the Retrieval Module, and 3) these two approaches are combined, where this approach optimizes the output of the Retrieval Module.
C.1.5 Latent feature space
Given the experience dataset $D$, context encoder $E$ decides the feature space that the Retrieval Module and the Video Plan Generator operate on. To implement $E$, we experiment with taking pre-trained vision encoders that are known to give good semantic representations, such as CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2024). To encode videos with these image encoders, we used a simple strategy that first encodes each frame independently and then concatenates the per-frame features, resulting in a single-vector video feature. A visualization of the constructed feature space is shown in Section D.1.
C.2 Rejection module
While the modules introduced in Section C.1 alone are sufficient to perform system identification essential for solving the tasks given a set of videos of interactions in the environment, the identified system parameters are often not perfect. This is especially the case when we only have a limited number of interactions with the environment. Thus, to effectively estimate the precise state of the environment, it is important that the system repeatedly actively interacts with the environment.
To actively interact with the environment, we use a video plan generator to generate plans for execution. To encourage the planner to explore novel plans and avoid being overly reliant on suboptimal beliefs, we employed a simple rejection-based sampling method. In the Rejection Module, the previously failed plans are stored in a data buffer. At each planning round, we generate $N$ plans $(v^{p}_{1},v^{p}_{2},\dots,v^{p}_{N})$ instead of one by conditioning the Video Plan Generator on $N$ independently sampled state embeddings from the Retrieval Module, yielding $N$ different plans with potentially different beliefs on environment parameter $\theta$. Let $\mathcal{F}$ be the set of previously failed plans stored in the data buffer, we first calculate the distance to the nearest failed plan for each plan $P_{i}$:
$d(v^{p}_{i},\mathcal{F})=\min_{F\in\mathcal{F}}K(v^{p}_{i},F)$
where $K$ is the distance function. Empirically, we found that L2 distance on raw pixel space works surprisingly well already.
We then select the plan $v^{p*}$ whose distance to its nearest failed plan is the largest (out of a set of generated video plans):
$v^{p*}=\arg\max_{v^{p}_{i}}d(v^{p}_{i},\mathcal{F})$
The selected $v^{p*}$ is later converted to low-level control signals through the Action Module.
The combination of Section C.1 and Section C.2 form the core of our Implicit State Estimation procedure. Interactions with the environment through the rejection module in Section C.2 obtain a buffer of interaction videos that help describe the implicit state of the environment. The retrieval module in Section C.1 then obtains an explicit latent that represents the implicit state of the environment.
C.3 Action module
Given a generated video plan, to convert the video into continuous actions to execute, we follow the concept of dense object-tracking from AVDC to recover actions from video plans. On certain interaction tasks, the object of interest is not applicable. For example, in a bar-picking task where the robot is required to grasp around the center of mass to succeed, knowing the future trajectory of the bar is not sufficient to decide the grasp location. In these tasks, we track the robot’s wrist instead. After obtaining the object/robot wrist trajectory from the tracking results, we convert the trajectory into actions using hand-crafted heuristic policies.
Appendix D Additional Experiment Results
D.1 State Embedding Feature Space
Figure 17 visualizes the feature space introduced in Section 4.1, using DINOv2 (Oquab et al., 2024) as the encoder. The plots are generated by applying t-SNE to sets of state embeddings from each environment, where for continuous modes we randomly sample four values for visualization. Our method yields a well-structured feature space in which different modes can be separated without additional data-processing, facilitating more effective downstream learning.
(a) Pick Bar (Continuous)
(b) Push Bar (Continuous)
(c) Slide Brick (Continuous)
(d) Open Box (Discrete)
(e) Turn Faucet (Discrete)
Figure 17: Feature space visualization. Top row: continuous parameter environments (Pick Bar, Push Bar, Slide Brick). Bottom row: discrete parameter environments (Open Box, Turn Faucet).
D.2 Training Data and Scaling Effects.
Table 8: Number of demonstrations used for each task. Each cell shows failed/successful demonstrations.
Push Bar Pick Bar Slide Brick Open Box Turn Faucet Total
Main exp (100%) 6000/250 6000/250 1560/130 20/20 20/20 13620/650
28% 1000/250 1000/250 1300/130 20/20 20/20 3360/650
The training data used in our main experiment is summarized in LABEL:tab:num_data, where the ”100%” row indicates the amount of data per environment. To account for the scarcity of data in real-world settings and to assess the effect of data scale on performance, we repeated the experiment with only 28% of the data. Results are presented in Table 9 and Figure 18.
Table 9: Scaling effect on performance. Reported metric is Replans until Success ($\downarrow$, lower is better). We compare performance with 28% vs. 100% data across each environment.
(a) Push Bar
Method 28% 100%
CQL 12.69 9.96
BC 11.24 10.26
AVDC 5.65 6.83
Ours 3.91 4.26
(b) Pick Bar
Method 28% 100%
CQL 11.57 9.22
BC 8.88 8.79
AVDC 3.60 4.41
Ours 3.56 3.84
(c) Slide Brick
Method 28% 100%
CQL 12.33 8.94
BC 7.34 9.24
AVDC 10.75 7.36
Ours 11.42 6.69
(d) Open Box
Method 28% 100%
CQL 3.70 3.32
BC 3.55 1.59
AVDC 1.84 1.82
Ours 1.39 1.25
(e) Turn Faucet
Method 28% 100%
CQL 4.68 5.63
BC 4.15 3.54
AVDC 1.78 1.67
Ours 1.59 1.39
(f) Overall (Normalized)
Method 28% 100%
CQL 3.36 2.56
BC 2.58 1.98
AVDC 1.32 1.31
Ours 1.16 1.00
28% Data100% Data112233CQLBCAVDCOurs
Figure 18: Overall scaling effect. Reported metric is Replans until Success ($\downarrow$ lower is better).
With 28% of the data, our method surpasses baselines in 4 out of 5 tasks and achieves the best overall performance by a substantial margin, and the aggregate results show a clear advantage. As more data becomes available, our approach improves further, scaling smoothly and consistently across tasks. Taken together, these results demonstrate that our method delivers consistent and robust gains, both under limited data and when scaling up.
D.3 Runtime analysis
We measured the average computation time over 100 iterations on a machine with an Intel Xeon W-2255 CPU and an NVIDIA RTX 3080 Ti GPU. Each iteration includes embedding retrieval, video generation (DDIM), plan selection (reject), and action decoding via optical flow. As shown in Table 10, planning takes only a few seconds per episode, making our framework practical for real-time interaction.
Table 10: Computation time (in seconds) averaged over 100 trials.
Component AVDC Ours Ours + Refine
Retrieve N/A 0.487 $\pm$ 0.008 11.876 $\pm$ 0.036
DDIM 0.777 $\pm$ 0.001 1.573 $\pm$ 0.004 1.599 $\pm$ 0.006
Reject N/A 1.276 $\pm$ 0.075 1.230 $\pm$ 0.068
Optical Flow 1.564 $\pm$ 0.021 1.661 $\pm$ 0.030 1.761 $\pm$ 0.028
Total 2.341 $\pm$ 0.021 4.997 $\pm$ 0.081 16.466 $\pm$ 0.082
Appendix E Limitation
To isolate the challenge of video-based belief refinement, our work assumes a reliable action module, attributing failures solely to planning errors. While this abstraction simplifies evaluation, it omits real-world execution noise, which future work could address through uncertainty-aware control. Our visually driven approach enables interpretability but struggles with visually ambiguous tasks (e.g., subtle directional differences); incorporating tactile or proprioceptive sensing could improve disambiguation. We also observe a “first-frame bias” in the video generator under low-data regimes, where plans become overly deterministic. Although we have shown in Section 5.6 that such bias can be mitigated via increasing the number of candidate plans $n$ and noise injection to the first-frame image, future methods may benefit from diversity-promoting strategies or explicit mode learning.